OpenStudy (anonymous):

solve the equations for the variable b)loga(5x)-loga (x+3)=1/3 loga 64 c) log2 x + log2 (x-2)-3=0

OpenStudy (lalaly):

\[\log_a \frac{5x}{x+3}= \log_a \sqrt[3]{64}\]\[\frac{5x}{x+3}=4\]